A tire placed on a balancing machine in a service station starts from rest and turns through 19.0 revolutions in 8.89 s before reaching its final angular speed. Calculate its angular acceleration.
Answer #1

using kinemati cs equations Angular acceleration, c here,9 19 0*2* π 2*38 *π (8.89) 38* π radians - 2 3.01 radIs
Answer #2

19*2*3.14=.5*a*8.89^2
solve for a

3.019 rad/s^2
Answer #3

s= 0.5* a t^2

19*2 pi= 0.5 alpha 8.89^2

angular acceleration alpha =3.02 rad/s^2